what is the anser for the following, 60% of the students in the computer course passed the test. if 25 students were randomly chosen at random.
1) what is the probabiliy that more than 18 students passed the test
2)what is the probability between 10 and 17 passed the test
3) calculate the expected mean
4)calculate the standard deviation
a soft drind bottling company wants to fill bottles labeled 360 ml such that bottles that contain less than 355.5 ml will be declared as non-quality standard and sold st substantial discont. the fills are normally distributed with a mean of 360 ml and standard deviation of 3.50ml. if 20,000 bottles are filled, approximatly HOW MANY will be declared as non-quality standard

Answer by stanbon(75887)

60% of the students in the computer course passed the test. if 25 students were randomly chosen at random.
----
Binomial Problem with n = 25 ; p = 0.6
----
1) what is the probabiliy that more than 18 students passed the test
P(19 <= x <= 25) = 1 - P(0<= x <=18) = 1-binomcdf(25,0.6,18) = 0.0736
-----------------------------------------------------------
2)what is the probability between 10 and 17 passed the test
binomcdf(25,0.6,17) - binomcdf(25,0.6,9) = 0.8333
-----------------------------------------------------------
3) calculate the expected mean = np = 25*0.6 = 15
--------------------
4)calculate the standard deviation = sqrt(npq) = sqrt(15*0.4) = 2.45
====================================================================
A soft drink bottling company wants to fill bottles labeled 360 ml such that bottles that contain less than 355.5 ml will be declared as non-quality standard and sold st substantial discount. The fills are normally distributed with a mean of 360 ml and standard deviation of 3.50ml. If 20,000 bottles are filled, approximatly HOW MANY will be declared as non-quality standard
----
Draw the picture of a normal curve
Put 360 in the middle and note that std = 3.50
Find the z value of 355.5:
z(355.5) = (355.5-360)/3.5 = -1.2857
--
Then P(z < -1.2857) = 0.0993
------------------------------------
If 20,000 bottles are filled, approximatly HOW MANY will be declared as non-quality standard
Ans: 0.0993*20000 = 1985
==============================
Cheers,
Stan H.
==============================